John and Peter shared a bunch of lynchees equally. John gave away 120 of his lychees and peter sold 744 of his lychees. Now John had 4 times as many lychees left as peter. How many lychees did Peter left?
Answer by lwsshak3(11628)

John and Peter shared a bunch of lynchees equally. John gave away 120 of his lychees and peter sold 744 of his lychees. Now John had 4 times as many lychees left as peter. How many lychees did Peter left?
**
let x=no. of lychees john and peter shared equally
x-120=no. of lychees john had left
x-744=no. of lychees peter had left
x-120=4(x-744)=4x-2976
3x=2856
x=952
x-744=208
no. of lychees peter had left=208